# Instruct-MusicGen: Unlocking Text-to-Music Editing for Music Language Models via Instruction Tuning

### v2

@@ 3.2.2 Text Fusion Module
First, the text is embedded and encoded by the T5 text encoder as $z_{\text{instruct}}=\text{T5}(X_{\text{instruct}})$. Let $W^{\prime}_{\rm Q}$, $W^{\prime}_{\rm K}$, and $W^{\prime}_{\rm V}$ denote the query, key, and value projection matrices of the pretrained cross-attention transformer layers respectively. Then, the query, key, and value of $z_{\text{instruct}}$ and $Z_{\text{music}}$ for the multi-modal cross-attention are calculated as follows:
$K_{\rm instruct}=z_{\text{instruct}}W^{\prime}_{\rm K},\hskip 14.22636ptQ_{\rm instruct}=Z_{\text{music}}W^{\prime}_{\rm Q},\hskip 14.22636ptV_{\rm instruct}=z_{\text{instruct}}W^{\prime}_{\rm V}.$ (9)
For efficient finetuning of the cross-attention module, we apply LoRA to the projection layers $W^{\prime}_{Q}$ and $W^{\prime}_{V}$. LoRA decomposes each of $W^{\prime}_{Q}$ and $W^{\prime}_{V}$ into two smaller low-rank matrices, significantly reducing the number of parameters requiring fine-tuning. The $Q_{\rm instruct}$ and $V_{\rm instruct}$ of cross-attention are then revised as follows:
$Q_{\rm instruct}=z_{\text{instruct}}(W^{\prime}_{\rm Q}+(W_{\rm Q}^{\text{A}})^{\mathsf{T}}W_{\rm Q}^{\text{B}}),$ (10)
$V_{\rm instruct}=z_{\text{instruct}}(W^{\prime}_{\rm V}+(W_{\rm V}^{\text{A}})^{\mathsf{T}}W_{\rm V}^{\text{B}}).$ (11)
During fine-tuning, $W^{\prime}_{Q}$, $W^{\prime}_{K}$, and $W^{\prime}_{V}$ are frozen, while $W^{\text{A}}_{Q}$, $W^{\text{B}}_{Q}$, $W^{\text{A}}_{V}$, and $W^{\text{B}}_{V}$ are trainable.
@@ 4 Experiments
4.1 Dataset
@@
    Slakh2100 dataset (Manilow et al., 2019). The Synthesized Lakh (Slakh) Dataset, originally derived from the Lakh MIDI Dataset v0.1, comprises high-quality audio tracks synthesised using sample-based virtual instruments. Known as Slakh2100, this dataset features 2100 tracks complete with corresponding MIDI files. These tracks are methodically divided into subsets designated for training (1500 tracks), validation (375 tracks), and testing (225 tracks), collectively amounting to approximately 145 hours of audio mixtures. This dataset primarily supports audio source separation tasks within a controlled, synthesised environment.
@@
    MoisesDB dataset (Pereira et al., 2023). In contrast to the Slakh2100, the MoisesDB dataset caters to source separation with a broader scope. It includes 240 real audio tracks sourced from 45 diverse artists spanning twelve musical genres. Uniquely, MoisesDB organises its tracks into a detailed two-level hierarchical taxonomy of stems, offering a varied number of stems per track, each annotated with textual descriptions. This setup facilitates a more comprehensive evaluation across different genres and stem configurations.
The rationale for selecting two datasets lies in their diverse configurations and common applications. While the Slakh dataset is traditionally utilized for training models tailored to a four-stem arrangement (e.g., Postolache et al., 2024), our model, Instruct-MusicGen, although initially trained on this dataset, is designed to generalise to various stem configurations. Conversely, models such as InstructME and AUDIT are trained on private or larger, more diverse datasets. By employing both Slakh2100 and MoisesDB, we ensure a comprehensive evaluation, allowing us to fairly compare the adaptability and performance of different models under varying conditions of data familiarity and complexity.
4.2 Data Preprocessing
We utilized the Slakh2100 dataset to construct an instruction-based dataset for our experiments, employing the following pipeline:
@@ 4.2 Data Preprocessing
    A data point was randomly selected from the Slakh training dataset.
@@
  3. 3.
  4. 4.
4.3 Experimental setup
For the finetuning of MusicGen, we jointly trained the audio fusion module and the text fusion module. The optimisation process utilised the AdamW optimiser, with a learning rate set at $5e^{-3}$. We use L2 loss over latent tokens as the training objective. Training incorporated a Cosine Annealing scheduler with an initial warmup of 100 steps. The training regimen extended over 5,000 steps with an accumulated batch size of 32, achieved through setting the batch size to 8 and using gradient accumulation over 4 iterations. The finetuning process was executed on a single NVIDIA A100 GPU and was completed within a span of two days.
4.4 Baselines
In this section, we explore various baseline models, each distinguished by their unique methodologies for handling audio data. We categorise these into three types: text-based music editing models, stem-wise modelling, and specialised models for single text-based tasks (see Table 1 for a comparison). Note we use "ME" to distinguish whether a model is a text-to-music editing model or not.
4.4.1 Text-based Music Editing Models
  1. 1.
    AUDIT (Wang et al., 2023): AUDIT is an instruction-guided audio editing model, consisting of a variational autoencoder (VAE) for converting input audio into a latent space representation, a T5 text encoder for processing edit instructions, and a diffusion network that performs the actual audio editing in the latent space. The system accepts mel-spectrograms of input audio and edit instructions, and generates the edited audio as output.
  2. 2.
    InstructME (Han et al., 2023): InstructME is a novel framework for instruction-guided music editing and remixing that leverages latent diffusion models. It enhances a U-Net with multi-scale aggregation to ensure consistency in the music before and after editing, and incorporates a chord progression matrix into the semantic space to improve melodic harmony. To handle extended musical pieces, InstructME utilizes a chunk transformer to recognize long-term temporal dependencies within music sequences.
  3. 3.
    M2UGen (Hussain et al., 2023): The M2UGen framework leverages large language models to comprehend and generate music across various modalities, integrating abilities from external models such as MusicGen (Copet et al., 2023) and AudioLDM 2 (Liu et al., 2023a). It is designed to stimulate creative outputs from diverse sources, showcasing robust performance in multi-modal music generation.
4.4.2 Specialised models
  1. 1.
    GMSDI (Postolache et al., 2024): GMSDI generalizes multi-source diffusion models to handle compositional music generation tasks without requiring pre-separated data. It allows for the flexible parameterization of an arbitrary number of sources and employs text embeddings to enhance semantic control over the music generation process. GMSDI uses the Slakh and Jamendo datasets for model training.
  2. 2.
    AudioSep (Liu et al., 2023c): AudioSep introduces the LASS-Net, which performs audio source separation based on natural language queries. This model effectively combines acoustic and linguistic data to isolate audio sources that match the language descriptions, significantly improving over conventional methods. AudioSep uses multiple datasets including AudioSet, VGGSound, AudioCaps, Clotho v2 and WavCaps dataset for model training.
Model ME Param size Dataset Hours (h) Training steps
@@ 4.4.2 Specialised models
Table 1: Comparison of different models.
4.5 Metrics
To comprehensively evaluate the quality and alignment of predicted music with ground truth and textual descriptions, we employ several metrics. The Fr’echet Audio Distance (FAD)(Kilgour et al., 2019) [^3] https://github.com/gudgud96/frechet-audio-distance. measures overall audio quality by comparing the distributions of feature embeddings between two sets of audio files. The CLAP score (CLAP)(Wu et al., 2023b) [^4] https://github.com/LAION-AI/CLAP. assesses how well the audio content aligns with corresponding textual descriptions, indicating semantic coherence. Kullback-Leibler Divergence (KL) [^5] https://github.com/haoheliu/audioldm_eval. evaluates the information loss by comparing the probability distributions of audio features from different sources, ensuring feature similarity between predicted and ground truth music. The Structural Similarity (SSIM)(Wang et al., 2004) metric, adapted from image quality assessment, evaluates the structural similarity between predicted music and the ground truth. Lastly, the Scale-Invariant Signal-to-Distortion Ratio (SI-SDR)(Roux et al., 2019) and Scale-Invariant Signal-to-Distortion Ratio improvement (SI-SDRi) (Isik et al., 2016) measure audio quality, particularly in source separation tasks, by quantifying the distortion relative to a reference signal and the improvement after processing, respectively. A detailed explaination can be found in Appendix A.
4.6 Results
Our evaluation of Instruct-MusicGen demonstrates its superior performance across various tasks compared to existing text-to-music editing baselines (AUDIT, InstructME, M2UGen) and specialized trained models (GMSDI, AudioSep). On the Slakh dataset (Table 2), Instruct-MusicGen excelled in adding, removing, and extracting stems, achieving the lowest Fréchet Audio Distance (FAD) and highest CLAP and SSIM scores in the addition task. It also significantly improved the signal-to-noise ratio (SI-SDR) in the removal task, showing balanced performance across all metrics and proving its robustness in various editing scenarios.  [^6] Since InstructME’s model weights and evaluation protocol are not public released, we only record the evaluation results in Table 5, which were reported in the original paper.
[…]
Add AUDIT 6.88 0.12 1.02 0.21 - -
InstructME* - - - - - -
M2UGen 7.24 0.22 0.99 0.20 - -
GMSDI 3.29 0.17 0.46 0.31 - -
Ours 3.75 0.23 0.67 0.26 - -
Remove AUDIT 15.48 0.07 2.75 0.35 -45.60 -47.28
InstructME* - - - - - -
M2UGen 8.26 0.09 1.59 0.23 -44.20 -46.13
GMSDI 1.30 0.12 0.23 0.78 5.67 3.74
AudioSep 0.90 0.18 0.63 0.65 1.38 -0.31
Ours 3.35 0.12 0.66 0.45 -2.09 -3.77
Extract AUDIT 15.08 0.06 2.38 0.42 -52.90 -50.16
InstructME* - - - - - -
M2UGen 8.14 0.11 2.15 0.31 -46.38 -43.53
GMSDI 2.91 0.12 0.28 0.74 2.73 5.58
AudioSep 2.49 0.15 1.00 0.62 -0.96 1.78
Ours 3.24 0.12 0.54 0.52 -9.00 -6.15
Similarly, in the MoisesDB dataset evaluations (Table 3), Instruct-MusicGen continued to demonstrate strong performance. It achieved competitive FAD scores and showed improvements in CLAP and SSIM metrics for both addition and removal tasks. While it did not always lead in SI-SDR, it consistently outperformed baseline models and was comparable to specialized models, highlighting its efficiency and effectiveness in text-to-music editing applications.
[…]
Add AUDIT 4.06 0.12 0.84 0.21 - -
InstructME* - - - - - -
M2UGen 5.00 0.18 0.83 0.20 - -
GMSDI 5.40 0.12 0.60 0.30 - -
Ours 3.79 0.18 0.35 0.35 - -
Remove AUDIT 10.72 0.10 2.46 0.34 -44.32 -57.10
InstructME* - - - - - -
M2UGen 3.75 0.13 1.27 0.19 -43.94 -56.73
GMSDI 1.42 0.09 0.41 0.65 -2.35 -15.13
AudioSep 1.25 0.13 0.49 0.68 4.97 -7.81
Ours 5.05 0.10 0.84 0.34 -13.70 -26.48
Extract AUDIT 6.67 0.07 1.97 0.45 -54.53 -56.17
InstructME* - - - - - -
M2UGen 5.74 0.08 1.91 0.25 -42.84 -44.49
GMSDI 4.95 0.08 1.32 0.54 -8.19 -9.83
AudioSep 1.81 0.14 1.01 0.65 1.14 -0.50
Ours 4.96 0.11 1.36 0.40 -21.39 -23.03
Despite these promising results, there are limitations to our approach. One significant limitation is the reliance on synthetic data from the Slakh2100 dataset for training, which may not fully capture the diversity and complexity of real-world music recordings. For example, while Instruct-MusicGen achieved a competitive FAD score of 3.35 and an SSIM of 0.45 in the removal task on the Slakh dataset, its SI-SDR improvement (SI-SDRi) was -3.77, indicating that while the audio quality is high, the separation quality could be enhanced. Similarly, in the extraction task, our model achieved a respectable FAD of 3.24 and SSIM of 0.52, but the SI-SDR and SI-SDRi metrics were -9.00 and -6.15, respectively, highlighting challenges in accurately isolating stems without introducing noise. These metrics suggest that while Instruct-MusicGen performs well in preserving audio fidelity and similarity, it struggles with signal clarity in more complex editing scenarios.
We provide further ablation studies in the Appendix A.  [^7] Demos are available at: https://bit.ly/instruct-musicgen.
5 Conclusion
In this paper, we introduced Instruct-MusicGen, a novel approach to text-to-music editing that fosters joint musical and textual controls. By finetuning the existing MusicGen model with instruction tuning, Instruct-MusicGen demonstrated its capability of editing music in various ways, including adding, separating and extracting a stem from music audio using textual queries, without the need for training specialized models from scratch. Also, it outperforms various baseline models that are dedicated to specific music editing tasks. Furthermore, our method uses significantly fewer resources than previous models, with a requirement of tuning only 8% of the parameters of the original MusicGen.
Instruct-MusicGen still has several limitations. Firstly, the generative editing tasks do not guarantee signal-level precision, as the process of predicting EnCodec tokens inherently involves some inaccuracy. Second, the finetuning process still relies on paired data, limiting the model’s applicability to more complex editing tasks where such data is not readily available. These constraints highlight the need for further research to improve signal accuracy and expand the model’s capabilities to handle a broader range of editing scenarios without heavily relying on paired datasets. Nonetheless, the ability to edit music with fewer computational resources opens up new avenues for creative expression and automation in music composition. Instruct-MusicGen represents a significant contribution to the field of text-to-music research, offering a robust and efficient tool for both musicians and researchers.
Appendix A Appendix / supplemental material
A.1 Metrics
@@
    Fréchet Audio Distance (FAD) [Kilgour et al., 2019] [^8] https://github.com/gudgud96/frechet-audio-distance. measures the similarity between two sets of audio files by comparing multivariate Gaussian distributions fitted to feature embeddings from the audio data. We use the FAD score to evaluate the overall audio quality of the predicted music.
@@
    CLAP score (CLAP) [Wu et al., 2023b] [^9] https://github.com/LAION-AI/CLAP. evaluates the alignment and relevance of audio content with corresponding textual descriptions, using models pretrained on both audio and language tasks to assess semantic coherence between audio and text. A high CLAP score indicates the predicted music aligns with the text description well.
@@
    Structural Similarity (SSIM) [Wang et al., 2004] is an image quality metric that we adapt to evaluate structural similarity between predicted music and ground truth.
@@
    Scale-Invariant Signal-to-Distortion Ratio (SI-SDR) [Roux et al., 2019] quantifies audio quality, especially in source separation tasks. It is scale-invariant, useful for varying audio volumes, and measures distortion relative to a reference signal. We use SI-SDR score to evaluate the signal loss of the predicted audio.
@@
    Scale-Invariant Signal-to-Distortion Ratio improvement (SI-SDRi) [Isik et al., 2016] extends SI-SDR, measuring the improvement in signal-to-distortion ratio before and after processing. It is commonly used in audio enhancement and separation contexts.
A.2 Ablations
A.2.1 Hyperparameter selection
The total trainable parameters include the input embedding layer, linear layers for joint embeddings learning, and gate scalars, and the linear layers has over 80% parameters in the total parameters. In this ablation we want to investigate whether the parameters can be less by revising the architecture of the linear layers.
The original linear layers has the shape of $(48,2048,2048)$, where 48 is the layer number of MusicGen model, and the $(2048,2048)$ is a linear transformation from embedding to layer-wise representation. We replace a single layer with an auto-encoder architecture, consisting of two linear layers: $(2048,M)$ and $(M,2048)$, where $M\ll 2048$.
We record and compare the loss on the validation set between the variant and the original instruct-MusicGen model, shown in Figure 3. The variation has $M=12$, which reduces the parameter size from 201M to 23M. From the experimental results, we find that even the paramater size becomes less, the instruction tuning still works for MusicGen, but the small parameter size limits the model’s representation ability, making the model performs worse than the original version.
Figure 3: The comparison between the original instruct-MusicGen and its variant with less parameters.
A.2.2 Comparison with original text encoder
We use an audio fusion module to enable MusicGen to accept additional audio input, allowing the model to process instruction text input without adding a new adapter. By finetuning the cross-attention module between the text encoder and the MusicGen transformer, we can fuse instructional knowledge into text information. In this ablation study, we examine the necessity of finetuning the text encoder. Since the text encoder is already capable of accepting text input, can it directly understand the semantic meaning of editing instructions?
To investigate this, we revised the architecture of Instruct-MusicGen’s text fusion module, configuring the music decoder to only accept cross-attention information from the original text encoder.
From Table 4, we observe that removing the text fusion module and relying on the original text encoder to interpret the instructions significantly decreases the model’s performance. Qualitative results from the predicted samples indicate that the original text encoder, which is designed for text description, struggles with the format of instructional commands. It fails to fully understand the semantics of keywords such as "extract," "remove," "no," and "only," which are not typically used to describe music but rather to indicate editing operations. Therefore, the presence of the text fusion module is crucial for MusicGen to correctly interpret the semantic meaning of instructions.
@@ A.2.2 Comparison with original text encoder
Add w/o. text fusion 5.26 0.15 0.70 0.25 - -
w. text fusion 3.75 0.23 0.67 0.26 - -
Remove w/o. text fusion 5.58 0.09 1.23 0.31 -10.32 -12.25
w. text fusion 3.35 0.12 0.66 0.45 -2.09 -3.77
Extract w/o. text fusion 7.58 0.04 1.87 0.42 -23.06 -20.21
w. text fusion 3.24 0.12 0.54 0.52 -9.00 -6.15
Table 4: Performance comparison between models with and without text fusion module across different tasks.
A.3 Experiments on other datasets
Since GMSDI only handle 4-stem music modelling, we construct a variant evaluation on Slakh dataset, where all stem can be considered, and the results are shown in Table 5.
@@ A.3 Experiments on other datasets
Add AUDIT 6.30 0.08 1.27 0.14 - -
InstructME* 3.15 - - - - -
M2UGen 5.80 0.21 0.88 0.17 - -
Ours 3.07 0.20 0.53 0.21 - -
Remove AUDIT 12.19 0.06 2.58 0.27 -45.40 -49.32
InstructME* 1.87 - - - - -
M2UGen 6.42 0.10 1.42 0.22 -42.72 -46.02
AudioSep 0.96 0.16 0.57 0.64 2.51 -0.78
Ours 2.95 0.12 0.64 0.39 -3.01 -6.31
Extract AUDIT 13.78 0.04 2.86 0.43 -54.95 -49.10
InstructME* 5.04 - - - - -
M2UGen 6.29 0.12 1.64 0.30 -45.65 -42.53
AudioSep 2.81 0.16 1.30 0.59 -2.22 0.91
Ours 3.25 0.08 0.85 0.45 -9.89 -6.77
Table 5: Comparison of Text-Based Music Editing Models on the Slakh dataset (full stems).
A.4 Social Impact
The introduction of text-based music editing models like Instruct-MusicGen has the potential to significantly impact the music industry by democratizing music creation and editing. By lowering the technical barriers to music editing, these models enable a broader range of individuals, including those without formal music training, to engage in creative processes. This can foster greater diversity in music production, as more people from various backgrounds can contribute their unique perspectives and ideas. Additionally, the efficiency and accessibility of text-based music editing can streamline workflows for professional musicians and producers, allowing them to experiment with different musical elements quickly and with less effort.
However, it is important to recognise the limitations and challenges associated with these advancements. The reliance on paired data for finetuning can restrict the model’s versatility in handling more complex editing tasks. Moreover, the inherent loss in generative editing might affect the precision of the output, potentially requiring additional finetuning by human experts. While these models offer new tools for creative expression, they are not a replacement for the nuanced understanding and expertise of skilled musicians. The social impact of text-based music editing models will largely depend on how they are integrated into existing workflows and their acceptance by the music community.
A.5 Audio samples
We provide audio samples in 4. From the results we can find the instruct-MusicGen performs the adding , removing and extracting tasks.

### v3

@@ 3.2.2 Text Fusion Module
The instruction is embedded and encoded by the T5 text encoder as $z^{\text{instruct}}=\text{T5}(X^{\text{instruct}})$. For efficient finetuning of the cross-attention module, we apply LoRA to the query and value projection layers. Thus, we expand Equation 10 as follows,
$Q_{l},K^{\text{instruct}}_{l},V^{\text{instruct}}_{l}=\text{QKV-Lora}(s^{\prime}_{l},z^{\text{instruct}}),$ (11)
$z^{\text{music}}_{l+1}=\text{CrossAttn}(Q_{l},K^{\text{instruct}}_{l},V^{\text{instruct}}_{l}).$ (12)
During fine-tuning, only query and value projection layers are trainable in the text fusion module.
@@ 4 Experiments
We conduct both subjective experiments and objective experiments for evaluation, and also provide example spectrograms in Figure 3.
4.1 Objective Experiments
4.1.1 Dataset
@@
    Slakh2100 dataset [18]. The Synthesized Lakh (Slakh) Dataset, originally derived from the Lakh MIDI Dataset v0.1, comprises audio tracks synthesised using high-quality sample-based virtual instruments. This dataset features 2100 tracks complete with corresponding MIDI files.
@@
    MoisesDB dataset [19]. The MoisesDB dataset includes 240 real audio tracks sourced from 45 diverse artists spanning twelve musical genres. Uniquely, MoisesDB organises its tracks into a detailed two-level hierarchical taxonomy of stems, offering a varied number of stems per track, each annotated with textual descriptions.
The rationale for selecting two datasets lies in their diverse configurations and common applications. While the Slakh dataset is traditionally utilised for training models tailored to a four-stem arrangement, our model, Instruct-MusicGen, although initially trained on this dataset, is designed to generalise to various stem configurations. Conversely, models such as InstructME and AUDIT are trained on private or larger, more diverse datasets. By employing both Slakh2100 and MoisesDB, we ensure a comprehensive evaluation, allowing us to fairly compare the adaptability and performance of different models under varying conditions of data familiarity and complexity.
4.1.2 Data Preprocessing
We utilised the Slakh2100 dataset to construct an instruction-based dataset for our experiments, employing the following pipeline:
  • A data point was randomly selected from the Slakh training dataset.
4.1.3 Experimental Setup
For the finetuning of MusicGen, we jointly trained the audio fusion module and the text fusion module. The optimisation process utilised the AdamW optimiser, with a learning rate set at $5\times 10^{-3}$. We use L2 loss over latent tokens as the training objective. Training incorporated a Cosine Annealing scheduler with an initial warmup of 100 steps. The training regimen extended over 5,000 steps with an accumulated batch size of 32, achieved through setting the batch size to 8 and using gradient accumulation over 4 iterations. The finetuning process was executed on a single NVIDIA A100 GPU and was completed within two days.
4.1.4 Baselines
In this section, we explore two baseline models, each distinguished by their unique methodologies for handling audio data.
@@ 4.1.4 Baselines
    AUDIT [13]: AUDIT is an instruction-guided audio editing model, consisting of a variational autoencoder (VAE) for converting input audio into a latent space representation, a T5 text encoder for processing edit instructions, and a diffusion network that performs the actual audio editing in the latent space. The system accepts mel-spectrograms of input audio and edit instructions, and generates the edited audio as output.
@@
    M2UGen [12]: The M2UGen framework leverages large language models to comprehend and generate music across various modalities, integrating abilities from external models such as MusicGen [2] and AudioLDM 2 [4]. It is designed to stimulate creative outputs from diverse sources, showcasing robust performance in multi-modal music generation.
Besides, InstructME can also perform instruction-guided music editing and remixing with latent diffusion models. We exclude it from comparison because InstructME’s model weights and evaluation protocol are not publicly released.
Model Param size Dataset Hours (h) Steps
@@
Table 1: Comparison of different models, where the param size numbers are trainable parameters and total parameters respectively. Our model has the lowest parameter size, and only requires 5K training steps.
4.1.5 Metrics
@@
    Fréchet Audio Distance (FAD) [38] [^4] https://github.com/gudgud96/frechet-audio-distance. measures the similarity between two sets of audio files by comparing multivariate Gaussian distributions fitted to feature embeddings from the audio data. We use the FAD score to evaluate the overall audio quality of the predicted music.
@@
    CLAP Score (CLAP) [39] [^5] https://github.com/LAION-AI/CLAP. is used in our experiments to measure the correspondence between the edited music and a target text. For the removal task, the target text is generated by deleting the name of the removed instrument from the original text.
@@
    Structural Similarity (SSIM) [40] is an image quality metric that we adapt to evaluate structural similarity between predicted music and ground truth.
@@
    Scale-Invariant Signal-to-Distortion Ratio (SI-SDR) [41] quantifies audio quality, especially in source separation tasks. It is scale-invariant, useful for varying audio volumes, and measures distortion relative to a reference signal. We use SI-SDR to evaluate the signal loss of the predicted audio.
@@
    Scale-Invariant Signal-to-Distortion Ratio improvement (SI-SDRi) [42] extends SI-SDR, measuring the improvement in signal-to-distortion ratio before and after processing. It is commonly used in audio enhancement and separation contexts.
To further investigate whether the model successfully adds, removes or extracts the instrument, we propose the P-Demucs score to evaluate the model performance. This metric specifically focuses on detecting the presence of a newly added instrument in the generated audio. It leverages the Demucs model, a source separation model, to isolate the target instrument from the audio. After separation, the root-mean-square energy (RMSE) of the isolated track is analyzed. For example, if the instruction is to "add guitar," a non-silent guitar track is regarded as a successful edit.
4.1.6 Objective Experiment Results
Our evaluation of Instruct-MusicGen demonstrates its superior performance across various tasks compared to existing text-to-music editing baselines (AUDIT, InstructME, M2UGen). On the Slakh dataset (Table 2), Instruct-MusicGen excelled in adding, removing, and extracting stems, achieving the lowest Fréchet Audio Distance (FAD) and highest CLAP and SSIM scores. It also significantly improved the signal-to-noise ratio (SI-SDR) in the removal task, showing balanced performance across all metrics and proving its robustness in various editing scenarios. Similarly, in the MoisesDB dataset evaluations (Table 3), Instruct-MusicGen demonstrated strong performance, with the best performance on most metrics over the three tasks.
@@ 4.1.6 Objective Experiment Results
Add AUDIT 6.88 0.12 1.02 0.21 0.53 - -
M2UGen 7.24 0.22 0.99 0.20 0.43 - -
Ours 3.75 0.23 0.67 0.26 0.80 - -
Remove AUDIT 15.48 0.07 2.75 0.35 0.33 -45.60 -47.28
M2UGen 8.26 0.09 1.59 0.23 0.70 -44.20 -46.13
Ours 3.35 0.12 0.66 0.45 0.76 -2.09 -3.77
Extract AUDIT 15.08 0.06 2.38 0.42 0.61 -52.90 -50.16
M2UGen 8.14 0.11 2.15 0.31 0.60 -46.38 -43.53
Ours 3.24 0.12 0.54 0.52 0.75 -9.00 -6.15
@@
Add AUDIT 4.06 0.12 0.84 0.21 0.50 - -
M2UGen 5.00 0.18 0.83 0.20 0.45 - -
Ours 3.79 0.18 0.35 0.35 0.77 - -
Remove AUDIT 10.72 0.10 2.46 0.34 0.41 -44.32 -57.10
M2UGen 3.75 0.13 1.27 0.19 0.72 -43.94 -56.73
Ours 5.05 0.10 0.84 0.34 0.78 -13.70 -26.48
Extract AUDIT 6.67 0.07 1.97 0.45 0.60 -54.53 -56.17
M2UGen 5.74 0.08 1.91 0.25 0.52 -42.84 -44.49
Ours 4.96 0.11 1.36 0.40 0.78 -21.39 -23.03
We find that all models exhibit negative SI-SDR and SI-SDRi scores, which is a common occurrence when evaluating generative models on a signal level. These metrics are typically designed for source separation tasks and are not entirely fair to generative models, as they penalise even minor discrepancies between the generated and original signals. Generative models like Instruct-MusicGen often focus on producing perceptually plausible audio rather than perfectly matching the original signal at a technical level.
4.2 Subjective Experiments
4.2.1 Experimental Setup
We conducted a subjective listening test to evaluate the model’s performance. [^7] This subjective test was approved by the ethics committee of Sony. This test involved disseminating an online survey within the Music Information Retrieval (MIR) community and our broader research network, which resulted in the collection of 30 complete responses. The gender distribution of the participants was 23 males (76.7%) and 7 females (23.3%). Regarding professional musical education experience, 4 participants (13.3%) had less than 1 year of experience, 13 (43.3%) had between 1 and 5 years, and 13 participants (43.3%) had more than 5 years of experience. For the data preparation, we randomly selected a subset of data points from the objective test dataset. Specifically, 6 audio samples were chosen, comprising 2 audio samples for each subtask (add, remove, extract). Each data point included results from the baseline models, our models, and the ground truth from the dataset.
4.2.2 Metrics
  1. 1.
    Instruction Adherence (IA) assesses how accurately the generated music follows the given editing instruction. In this experiment, participants rate the generated music on a scale from 1 to 5, where 1 indicates that the instruction was not followed at all, and 5 indicates that the instruction was followed perfectly. For example, if the instruction is "Remove Drums," a rating of 1 would mean that the drums were not removed at all, while a rating of 5 would mean that the drums were completely removed.
  2. 2.
    Audio Quality (AQ) evaluates the overall audio quality of the generated music in comparison to the original music. Participants rate the audio quality on a scale from 1 to 5, where 1 represents very poor quality with significant degradation compared to the original music, and 5 represents excellent quality, as good as or better than the original music. This metric helps in understanding how the editing process affects the overall sound quality of the music.